Solve the following congruences x≡1 mod3, x≡2mod4


x≡3mod5

1
Expert's answer
2021-12-14T01:20:02-0500

Let us solve the following system of congruences:


"\begin{cases}
x≡1 \mod 3\\
x≡2\mod 4\\
x≡3\mod 5
\end{cases}."


The first congruence is equivalent to the equality "x=1+3t,\ t\in\Z." Let us put this in the second conqruence. Then we have "1+3t\equiv 2 \mod 4." The last conqruence is equivalent to "-3t\equiv -1 \mod 4," and hence to "(4-3)t\equiv (4-1) \mod 4," that is "t\equiv 3 \mod 4." It follows that "t=3+4s," and hence "x=1+3(3+4s)=10+12s." Let us put this in the last conqruence. Then we get "10+12s≡3\mod 5," which is equivalent to "12s≡-7\mod 5," and hence to "2s≡-2\mod 5." It follows that "s\equiv-1\mod 5," and hence "s=-1+5k." We conclude that "x=10+12s=10+12(-1+5k)=-2+60k." Therefore, the solution of the system is "x\equiv -2\mod 60" or "[-2]_{60}."